OpenStudy (anonymous):

Find the amplitude and period of the function? -5sin(4pix)

OpenStudy (anonymous):

So a general formula for a periodic function is as follows: \[ A\sin{2\pi f x} \] Where \(|A|\) will be the amplitude, and \(f\) will be the frequency, which is one over the period (\(P=\frac{1}{f}\)).

OpenStudy (anonymous):

Can you get the answer from that?

OpenStudy (anonymous):

I don't understand the P=1/f part?

OpenStudy (anonymous):

Well, P is the period. And \(f\) is the frequency. That's just saying the reciprocal of the frequency is the period.

OpenStudy (anonymous):

In the argument of your function you have 4*Pi*x. Your objective should be to make this match the form of 2*Pi*f*x. Then, your F is the denominator of your P=1/f.

OpenStudy (anonymous):

Ok, Thanks. One more y= 7sin (1/2(x+pi/3))

OpenStudy (anonymous):

Well, a more general form of the equation is: \[ A\sin(2\pi f x + \phi) \] This means all you have to do is multiply out the things inside the parentheses, and then fit them to find what your \(f\) is.